x+2 = cube root(12x+8)
thanks is advance

Answer by ankor@dixie-net.com(22740)

x+2 = cube root(12x+8)
:
Cube both sides:
(x+2)(x+2)(x+2) = 12x + 8
:
(x+2)*(x^2 + 4x + 4) = 12x + 8
:
x^3 + 6x^2 + 12x + 8 = 12x + 8
:
x^3 + 6x^2 + 12x - 12x + 8 - 8 = 0; arrange all terms on the left
:
x^3 + 6x^2 = 0
:
x^2(x + 6) = 0; factored out x^2
:
x = 0
and
x = -6
:
:
Check both solutions:
x = 0
0 + 2 = cubeRt(0 + 8)
2 = cubeRt(8)
and
-6 + 2 = cubeRt(-72 + 8)
-4 = cubeRt(-64)